OpenStudy (anonymous):

integrate tan^3 xsec x dx...!

OpenStudy (anonymous):

\[\int\limits \tan^3 x secx dx\]

OpenStudy (anonymous):

what to do?

OpenStudy (experimentx):

\[ \tan^3 x \sec x = \tan x \sec x ( \sec^2 x - 1)\] seems like it will get the same

OpenStudy (anonymous):

yeah ..next step

OpenStudy (anonymous):

note that d(sec x) =sec x tan x dx

OpenStudy (anonymous):

yes all right

OpenStudy (anonymous):

what's next..

OpenStudy (anonymous):

let u = sec x then proceed. du= sec x tan x dx so it all becomes \(\int (u^2 - 1) du\)

OpenStudy (anonymous):

dear are you solving integral by parts?

OpenStudy (anonymous):

or by substituion?

OpenStudy (anonymous):

it will become like this i guess u^3/3 - u +c

OpenStudy (anonymous):

\( \int (\sec^2-1) \tan x \sec x dx\) =\( \int (\sec^2-1) d(\sec x)\) =\( \int (u^2-1) du\), where u=\sec x\) by substitution.

OpenStudy (anonymous):

yes next step..

OpenStudy (anonymous):

eh? \(\int u^n du= \frac{u^{n+1}}{n+1} +C\)

OpenStudy (anonymous):

after applying we ll get as i above wrote ?

OpenStudy (anonymous):

you'll get the answer....from integrating the above eq.

OpenStudy (anonymous):

wait let me write what i am getting.

OpenStudy (anonymous):

\[\frac{ \sec^3x }{ 3 }- secx +c = answer ???\]

OpenStudy (anonymous):

it should be, yeah. :) you can always check your answers with wolfram.

OpenStudy (anonymous):

wolfram.?

OpenStudy (anonymous):

but answer is differnet in my book?

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }(secxtan^2x-2secx)+c =answer ( book)\]

OpenStudy (anonymous):

@Shadowys

OpenStudy (anonymous):

just take sec x out. and also, the integral, \(\int du=u\)

OpenStudy (anonymous):

did't get..

OpenStudy (anonymous):

\(\int (u^2-1) du = \int u^2 du - \int du\) following the integral, it becomes, \(\frac{u^3}{3} - u+C\)

OpenStudy (anonymous):

i have already written this.Scroll up a bit

OpenStudy (anonymous):

i just wanna get the same answer as in my book...:)

OpenStudy (anonymous):

oh. unfortunately, either your book's wrong, or the computer integration is wrong. because wolfram agrees with you....lol

OpenStudy (anonymous):

yeah i have seen there but what you think about book's anwer is that wrong,there is really less possibilit though and yes lolz?

OpenStudy (anonymous):

i have three books with same answer(:

OpenStudy (anonymous):

lol i dun think wolfram's wrong, in anycase....lol

OpenStudy (anonymous):

if we solve this by ''integration by parts'' rather tha by substituion?

OpenStudy (anonymous):

\[\int\limits uvdx= u \int\limits vdx - \int\limits(\int\limits vdx) u'dx\]

OpenStudy (anonymous):

same way, but this time i still let du=d sec x tan x dx. and v=sec^2 -1 sorry, gtg, but i think you'll still get the same answer...(or that's what my calc tells me)

OpenStudy (anonymous):

@Shadowys okay i think my all books are wrong this time.(there is no benifit or edge having more than one books) :)!

OpenStudy (anonymous):

thanks i took your time